OpenStudy (anonymous):

1/sin^2 - 1/tan^2 = 1

OpenStudy (mathstudent55):

What are you trying to do?

OpenStudy (anonymous):

Its the things where you have to use the identities.

OpenStudy (mathstudent55):

Ok. You want to prove the identity.

OpenStudy (mathstudent55):

Multiply both sides by sin^2 theta

OpenStudy (mathstudent55):

\(\dfrac{1}{\sin^2 \theta} - \dfrac{1}{\tan^2 \theta} = 1\) \(\dfrac{1}{\sin^2 \theta} \times \sin^2 \theta - \dfrac{1}{\tan^2 \theta} \times \sin^2 \theta= 1 \times \sin^2 \theta\) Now use the identity \(\tan \theta = \dfrac{\sin \theta}{\cos \theta} \) on the left side to deal with \(\tan^2 \theta \).